OpenStudy (fernandez555):

Does the point you select matter when your write a point-slope equation? explain. why it is important to understand any limitations on the domain and range?

OpenStudy (pratyush5):

According to how i have learned it, in point slope form we are given only one point and slope. The question where two points are given is called two point form. It is solved by y - y1 = (x- x1) (y2 - y1/x2-x1). This is how i was taught .

OpenStudy (fernandez555):

@kelliegirl33 can you give me the equation in point slope form for points a (3,4) and point b (5,5) ???

OpenStudy (anonymous):

lets check.. two points (1,2) and (3,4) slope(m) = (4 - 2) / (3 - 1) slope(m) = 2/2 = 1 point slope : y - y1 = m(x - x1) slope(m) = 1 (1,2) x1 = 1 and y1 = 2 now we sub y - 2 = 1(x - 1) point slope form : y - y1 = m(x - x1) slope(m) = 1 (3,4) x1 = 3 and y1 = 4 now we sub y - 4 = 1(x - 3) as you can see, they are different point slope forms, but when you turn them into slope intercept forms... y - 2 = 1(x - 1) y - 2 = x - 1 y = x - 1 + 2 y = x + 1 and y - 4 = 1(x - 3) y - 4 = x - 3 y = x - 3 + 4 y = x + 1 the answers will come out the same. So, no, it does not matter which point you choose

OpenStudy (fernandez555):

@kelliegirl33 with steps if possible

OpenStudy (anonymous):

find the slope first.. slope(m) = (y2 - y1) / (x2 - x1) (3,4) x1 = 3 and y1 = 4 (5,5) x2 = 5 and y2 = 5 now we sub slope(m) = (5 - 4) / (5 - 3) slope(m) = 1/2 y - y1 = m(x - x1) (5,5) x1 = 5 and y1 = 5 slope(m) = 1/2 time to sub y - 5 = 1/2(x - 5) ==> point slope form

OpenStudy (fernandez555):

@kelliegirl33 last question. can you convert each of these equations to slope intercept form? 1. y - 4 = ½ x - 3 2. y - 5 = 1/2(x - 5)

OpenStudy (anonymous):

1. y - 4 = 1/2(x - 3) -- distribute through the parenthesis y - 4 = 1/2x - 3/2 -- add 4 to both sides y = 1/2x - 3/2 + 4 -- common denominator is 2 y = 1/2x - 3/2 + 8/2 -- combine like terms y = 1/2x + 5/2 2. y - 5 = 1/2(x - 5) -- distribute y - 5 = 1/2x - 5/2 -- add 5 to both sides y = 1/2x - 5/2 + 5 -- common denominator is 2 y = 1/2x - 5/2 + 10/2 -- combine like terms y = 1/2x + 5/2